# MaskInversion: Localized Embeddings via Optimization of Explainability Maps

### v1

@@ 3.2 Localized Embedding Learning via Explainability Map Optimization
The inputs to our method are an image $x\in\mathbb{R}^{3\times W\times H}$ and a binary query mask $\textbf{m}=(m_{i,j})\in\mathbb{R}^{W\times H}$, $m_{i,j}\in\{0,1\}$, specifying a region of interest. Our objective is to derive a localized embedding token $LET_{\textbf{m}}\in\mathbb{R}^{d}$ that generates an explainability map that corresponds to the masked region.
Embedding Token Initialization . We initialize the localized embedding token $LET_{\textbf{m}}^{(0)}$ by copying the global [CLS] token produced by the foundation model, $LET_{\textbf{m}}^{(0)}=z^{0}\in\mathbb{R}^{d}$. We then compute the cosine similarity between the embedding token and the average of the [CLS] and all patch tokens following (Bousselham et al., 2024) as the activation score for the explainability map:
$s^{(0)}=\textbf{cos}\left(LET_{\textbf{m}}^{(0)},\bar{\textbf{z}}\right)\in\mathbb{R},$ (2)
where $\bar{\textbf{z}}=\frac{1}{n}\sum_{p}z_{p}$ represents the patch and [CLS] token of the ViT averaged across the spatial dimensions, and cos denotes the cosine similarity. Following the process described in Section 3.1, the score is used to compute the explainability map denoted as $\mathbf{E}^{(0)}=E(s^{(0)})\in\mathbb{R}^{W\times H}$, with each element $\mathbf{E}^{(0)}_{i,j}\in[0,1]$. This map $\mathbf{E}^{(0)}$ indicates the regions within the image that the initial embedding $LET_{\textbf{m}}^{(0)}$ predominantly focuses on. Since the localized embedding is initialized with the [CLS] token our initial explainability map corresponds to the explainability map of the [CLS] token.
Embedding Token Optimization. To refine the initial guess and guide the embedding token representation towards the query mask, we treat the mask localized embedding $LET_{\textbf{m}}\in\mathbb{R}^{d}$, corresponding to the query mask $\mathbf{m}$, as a learnable vector with $d$ parameters. We supervise the learning of this vector by optimizing its parameters for $K$ steps, with $k\in\{0,..,K\}$ so that, for each optimization step $k$, the resulting explainability map $\mathbf{E}^{(k)}$ for this token resembles the query mask m. We achieved this through iterative gradient descent. Specifically, we quantify the discrepancy between the explainability map and the query mask using a soft Dice loss, as commonly employed in segmentation tasks Milletari et al. (2016); Cheng et al. (2021); Bousselham et al. (2022) measuring region similarity:
@@
where $\mathrm{intersection}(\mathbf{E}^{(k)},\textbf{m})$ and $\mathrm{union}(\mathbf{E}^{(k)},\textbf{m})$ are the intersection, realized by elementwise multiplications, and union, realized by elementwise addition, of the explainability map and the binary mask, respectively, and $\epsilon$ is a small constant to avoid division by zero. The Dice loss is minimized by optimizing the localized embedding $LET_{\textbf{m}}$ parameters over $K$ iterations of gradient descent to yield the final embedding $LET_{\textbf{m}}=LET_{\textbf{m}}^{(K)}$.
Regularization Loss. The method, as described so far, will capture the representation of the indicated region. This can lead to the effect that the final representation $LET_{\textbf{m}}$ is less aligned with the image itself, thus discarding any image information. But it can sometimes be helpful to have both a good region representation together with general image context. We, therefore, propose an add-on auxiliary regularization loss that forces the localized token embedding $LET_{\textbf{m}}^{(k)}$ at each step $k$ to remain within the manifold of the image encoder:
@@
where $\alpha\in\mathbb{R}$ is a hyperparameter that modulates the influence of the regularization loss. It allows us to regulate how much region vs. global information should be encoded in the output token embedding. We found that this sepcifically helps for tasks that need context knowledge such as referring expressions, while ‘object-only’ tasks such as region-based/localized classification do not profit from such an alignment, thus setting $\alpha=0$ for those cases.
3.3 Faster mask inversion via gradient decomposition
The derivation of the explainability map necessitates the calculation of a gradient, and similarly, each gradient descent iteration requires the computation of a gradient with respect to the loss function $\mathcal{L}$. Consequently, this iterative process requires the evaluation of second-order derivatives of the form $\frac{\partial\mathcal{L}}{\partial LET_{\textbf{m}}^{(k)}}(LET_{\textbf{m}}^{(k)},\nabla\textbf{A})$, which can be computationally intensive and numerically unstable.
To enhance the computational efficiency of this process, it is advantageous to obviate the need for backpropagation to generate explainability maps at each iteration. We propose a gradient decomposition strategy that simplifies the gradient computation associated with the explainability method. For a given iteration $k$, the gradient decomposition can be expressed as follows:
@@ 4.1 Downstream Tasks
In the following, we will give a brief overview of these tasks, their metrics and datasets. Please see the Appendix B for all details.
@@
Referring Expressions To assess the proposed method’s ability to capture localized properties, we evaluate it for referring expression classification. Given an image and a set of masks, we generate an embedding for each mask within an image and match the generated region embeddings to a set of text queries (referring expressions) encoded with the respective text encoder. The query mask whose localized embedding exhibits the highest cosine similarity with the text embedding is selected. We employ standard referring expression datasets: PhraseCut Wu et al. (2020), RefCOCO, and RefCOCO+ Kazemzadeh et al. (2014), reporting top-1, top-5, top-10 accuracy, mean Intersection over Union (mIoU) and overall Intersection over Union (oIoU).
@@
Class Retrieval Zero-shot classification requires classifying an image by matching its visual embedding with the textual description of the classes present in the dataset. Here, we propose to increase the granularity by using it to classify a specific region of the image: given a query mask of an object, classify it by matching its localized embedding to the text embeddings of the classes in the datasets. For this, we leverage two semantic segmentation datasets, PascalVOC Everingham et al. (2015) and PascalContext Mottaghi et al. (2014), and one instance segmentation dataset, MSCOCO Lin et al. (2014). The performance is evaluated using the top-1, top-5, and top-10 accuracy. Finally, we challenge the proposed method in a large-scale open-vocabulary setting. We utilize a subset of the OpenImagesV7 Benenson & Ferrari (2022), which offers mask annotations for a diverse array of objects across 350 unique classes.
@@
Localized Captioning Traditionally, image captioning models generate captions for entire images based on the visual representation provided by an image encoder. In contrast, we aim to evaluate our method’s ability to focus the captioner on a specific image region while maintaining contextual relevance. To this end, we leverage a pretrained image captioner, CLIPCap Mokady et al. (2021), and provide it with the localized embedding token of a query mask to generate a caption. CLIPCap is trained on top of the CLIP vision encoder and feeds its [CLS] token to GPT-2Radford et al. (2019) to produce a caption. Here, we feed the localized embeddings of MaskInversion as a drop-in replacement of the CLIP [CLS] token to the captioner without any finetuning. As no dataset directly supports this evaluation type, we adapted an existing dataset, PhraseCut. To quantitatively evaluate the generated localized captions, we match the generated caption to the set of ground truth referring expressions for this image using the text encoder from CLIP (ViT-L/14 by OpenAI). We consider the caption correct if the cosine similarity between the generated caption and the ground truth referring expression for this mask is the highest. The reported metric for this task is the top-1 accuracy.
@@
Implementation Details The proposed method is evaluated using pretrained CLIP vision-language models. For ViT-B/32, ViT-B/16, and ViT-L/14, we used the original weight from OpenAI Radford et al. (2021), and for ViT-H/14, we used the weights "laion2b_s32b_b79k" from the OpenCLIP library Cherti et al. (2023); Schuhmann et al. (2022). For the MaskInversion process, we use AdamW optimizerKingma (2014) with $10$ gradient descent iterations. For the loss equation 5, we set $\alpha$ to $5$ for RefCOCO and RefCOCO+, and to 00 for all other datasets.
@@
RedCircleShtedritski et al. (2023) ViT-B/16 21.5 72.3 90.3 42.5 43.2 32.7 42.5 43.3 33.5
FGVPYang et al. (2024) ViT-B/16 35.9 83.5 95.2 42.6 43.2 33.3 48.0 48.7 38.0
AlphaCLIPSun et al. (2024) ViT-B/16 34.0 80.0 93.6 43.4 44.0 38.1 44.2 44.7 39.7
@@
[…]
4.2 Comparison to State-of-the-Art
Referring Expression Retrieval Table 1 presents the results on referring expression datasets. For related approaches, as there is no directly comparable setting, we provide both, reported as well as reproduced results. Note that the original evaluation settings can vary for different methods. For reproduced results, indicated by *, we adapt the evaluation setting to the case where ground truth masks are used as described in Sec. 4.1. We used the code provided by the authors of each method, forward each image together with the groundtruth masks of MSCOCO, and match the resulting representation to the text embedding of the respective backbone. We further compare with the following baselines: CLIP refers to the general CLIP baseline by using the image CLS token, Crop uses the CLS token of cropped region by forwarding only this region through CLIP, and Masked Crop refers to forwarding the full image, but keeping only the masked region and replacing all other pixels with the average pixel value of the dataset. For PhraseCut, MaskInversion outperforms all entertained baselines, irrespective of the model size. On RefCOCO and RefCOCO+, MaskInversion also achieves SOTA performance. Interestringly, MaskInversion performance scall-well when the backbone size increases, establishing new SOTA on every dataset when ViT-H/14 is used.
Class Retrieval Analysis Table 2 compares MaskInversion to other methods for the case of zero-shot class retrieval, keeping the same setting as detailed under Referring Expression Retrieval. MaskInversion again performs well compared to other methods on semantic segmentation datasets, such as PascalVOC and PascalContext.Furthermore, MaskInversion also exhibits good performance on the instance segmentation dataset COCO. These results demonstrate that MaskInversion can effectively direct the attention of the foundation model to multiple instances of the same object class at the same time, as well as to a single instance. Here, MaskInversion also outperforms the recently proposed AlphaCLIP Sun et al. (2024), which fine-tunes CLIP with millions of mask-text pairs annotations, thereby demonstrating its ability to excel without the need to fine-tune CLIP. Finally, looking at the results on OpenImagesV7, which features a significantly larger vocabulary of 350 classes, we can see that methods like AlphaCLIP, which are specifically trained for such tasks, perform well. However, MaskInversion still outperforms all other methods we compared, demonstrating its capability to handle large vocabularies.
@@ 4.2 Comparison to State-of-the-Art
RedCircle*Subramanian et al. (2022) 47.5 92.9 97.7 21.3 45.0 57.4 28.8 63.0 77.3 40.5 75.8 84.5
AlphaCLIP*Sun et al. (2024) 52.6 85.9 93.8 27.7 60.9 75.1 30.9 55.9 70.3 43.0 77.4 84.3
FGVP*Yang et al. (2024) 71.8 93.6 98.3 32.6 58.9 72.4 35.9 62.2 72.6 39.4 75.6 84.6
RIS*Yu et al. (2023) 78.0 95.2 98.1 38.1 62.7 74.3 43.6 65.3 72.4 34.5 66.5 75.8
B/32 MaskInversion 79.5 96.4 8.9 46.7 74.9 84.6 38.0 65.8 78.4 42.6 78.8 86.6
B/16 MaskInversion 85.4 96.4 98.8 58.1 83.7 90.5 44.7 71.61 83.0 46.3 80.4 87.9
@@
Figure 3: Localized Embedding Visualizations: visualisation of the learned localized embedding using (left) a pretrained diffusion model; (right) an image captioner. In both cases, all the models are kept frozen and only the global feature representation of the vision encoder is replaced by the output of MaskInversion depending on the query mask.
4.3 Localized Captioning Analysis
Localized captions We further consider the performance of MaskInversion against CLIP and AlphaCLIP for localized captioning in Table 3. We use CLIPCap Mokady et al. (2021) as the captioner and replace the CLIP image encoder with either AlphaCLIP or the output of MaskInversion without any fine-tuning. We observe that MaskInversion demonstrates the ability to focus the captioner on the area of interest, as the accuracy is more than doubled when using MaskInversion versus only using CLIP.
Method Acc
CLIP 20.1
AlphaCLIP 31.8
MaskInversion 48.4
Table 3: Localized captioning: Given a query mask, the goal is to generate a caption that corresponds to the region highlighted by the mask. CLIPCap is used to generate the caption with CLIP-ViT-B/16.
Moreover, MaskInversion also significantly AlphaCLIP despite not involving any fine-tuning of the CLIP model.
Qualitative Results Figure 3 presents further qualitative examples of the localized captions generated by MaskInversion+CLIPCap for different query masks. These visualizations complement the quantitative benchmarks. The proposed method demonstrates a high degree of precision in focusing the captioning on specific image regions dictated by the query masks, as e.g. water and first are in are highly separated.
4.4 Mask Embedding for Image Diffusion
To further visualize the concepts captured in the learned representation output by MaskInversion, we employed $\lambda$-ECLIPSE Patel et al. (2024), a state-of-the-art diffusion model. This model accepts a visual embedding from a ViT-bigG/14 CLIP model along with a text prompt, producing variations of the input image that correspond to the prompt. Utilizing the default settings of $\lambda$-ECLIPSE as described in Patel et al. (2024), we conducted several experiments to generate images based on different query masks used for the MaskInversion process.
4.5 Ablations
@@
Erosion 42.7
Dilation 44.3
Mask 44.7
Table 4: Mask Quality Ablation: assessment of the mask quality impact on MSCOCO for the Class Retrieval task.
Figure 3 illustrates how the generated images vary depending on the mask used. This variation shows the effectiveness of MaskInversion in producing localized and contextualized embeddings. The images clearly focus on the objects or groups of objects within the bounds of the query mask, confirming that MaskInversion directs the model’s attention to specific parts of the image. Moreover, we observe that the final explainability map generated by LeGradBousselham et al. (2024) is focused on the area covered by the query mask, validating the effectiveness of the proposed optimization process.
##\##Mask Decomp. Sec.$\downarrow$
@@
Table 5: Gradient Decomposition Ablation: Runtime using or not using gradient decomposition as described Sec.3.3 (ViT-B/16) for different number of masks, for 10 gradient descent steps.

### v2

@@ 3.2 Localized Embedding Learning via Explainability Map Optimization
We denote the input for the proposed method as an image $x\in\mathbb{R}^{3\times W\times H}$ and a binary query mask $\textbf{m}=(m_{i,j})\in\mathbb{R}^{W\times H}$, $m_{i,j}\in\{0,1\}$, specifying a region of interest. Our objective is to derive a localized embedding token $LET_{\textbf{m}}\in\mathbb{R}^{d}$ that generates an explainability map that corresponds to the masked region.
Embedding Token Initialization. We initialize the localized embedding token $LET_{\textbf{m}}^{(0)}$ by copying the global [CLS] token produced by the foundation model, $LET_{\textbf{m}}^{(0)}=z^{0}\in\mathbb{R}^{d}$ (Step 0 in Fig.2). We then compute the cosine similarity between the embedding token and the average of the [CLS] and all patch tokens as the activation score for the explainability map (Step $1$ in Fig.2)
$s^{(0)}=\mathrm{cos}\left(LET_{\textbf{m}}^{(0)},\bar{\textbf{z}}\right)\in\mathbb{R},$ (1)
where $\bar{\textbf{z}}=\frac{1}{n}\sum_{p}z_{p}$ represents the combined patch and [CLS] token representation averaged across the spatial dimensions, and cos denotes the cosine similarity. The resulting similarity score is used to compute the explainability map denoted as $\mathbf{E}^{(0)}=E(s^{(0)})\in\mathbb{R}^{W\times H}$, with each element $\mathbf{E}^{(0)}_{i,j}\in[0,1]$. This map $\mathbf{E}^{(0)}$ indicates the regions within the image that the initial embedding $LET_{\textbf{m}}^{(0)}$ predominantly focuses on. Since the localized embedding is initialized with the [CLS] token our initial explainability map corresponds to the explainability map of the [CLS] token.
Embedding Token Optimization. To refine the initial estimate and guide the embedding token representation towards the query mask, we treat the mask localized embedding $LET_{\textbf{m}}\in\mathbb{R}^{d}$, corresponding to the query mask $\mathbf{m}$, as a learnable vector with $d$ parameters. We supervise the learning of this vector by optimizing its parameters so that the resulting explainability map $\mathbf{E}^{(k)}$ for this token resembles the query mask m.
We achieve this goal via iterative gradient descent. Specifically, we quantify the discrepancy between the explainability map and the query mask using a soft Dice loss, as commonly employed in segmentation tasks (Milletari et al., 2016; Cheng et al., 2021) for measuring region similarity:
@@
where $\mathrm{intersection}(\mathbf{E}^{(k)},\textbf{m})$ and $\mathrm{union}(\mathbf{E}^{(k)},\textbf{m})$ are the intersection, realized by elementwise multiplications, and union, by elementwise addition, of the explainability map and the binary mask; $\epsilon$ is a small constant to avoid division by zero. We minimize the Dice loss by optimizing the localized embedding $LET_{\textbf{m}}$ parameters over $K$ iterations of gradient descent to achive the final embedding $LET_{\textbf{m}}=LET_{\textbf{m}}^{(K)}$ (see Step $K$ in Fig.2).
Handling Multiple and Overlapping Masks. It is worth noting that the optimization process described above is instantiated independently for each query mask. Consequently, MaskInversion naturally handles images containing dense object clusters or overlapping masks without cross-mask interference. Since the Dice loss (Eq. 2) is computed solely between the generated explainability map and the specific binary query mask, the presence of other objects or overlapping regions does not affect the convergence or quality of the target embedding.
Regularization Loss. The method as described so far will capture the representation of the region indicated by the query mask. This can cause the final representation $LET_{\textbf{m}}$ to be less aligned with the global context of the overall image. However, it can be helpful to have both a good region representation together with global image context. We therefore introduce an add-on auxiliary regularization loss that forces the localized token embedding $LET_{\textbf{m}}^{(k)}$ to remain close to the original image embedding by minimizing the respective distance:
@@
where $\alpha\in\mathbb{R}$ is a hyperparameter modulating the influence of the regularization loss. It allows us to regulate how much regional vs. global information should be encoded in the output token embedding. We found that this helps for tasks that need context knowledge such as referring expressions retrieval(Wu et al., 2020), while ‘object-only’ tasks such as localized classification do not profit from such an alignment.
Faster mask inversion via gradient decomposition The derivation of the explainability map necessitates the calculation of a gradient, and similarly, each gradient descent iteration requires the computation of a gradient with respect to the loss function $\mathcal{L}$. Consequently, this iterative process requires the evaluation of second-order derivatives of the form $\frac{\partial\mathcal{L}}{\partial LET_{\textbf{m}}^{(k)}}(LET_{\textbf{m}}^{(k)},\nabla\textbf{A})$.
These can be computationally intensive and numerically unstable. To enhance the computational efficiency of this process, it is advantageous to obviate the need for backpropagation to generate explainability maps at each iteration. We propose a gradient decomposition strategy that simplifies the gradient computation associated with the explainability method. For a given iteration $k$, the gradient decomposition can be expressed as follows:
@@ 4.1 Downstream Tasks
We give a brief overview of the evaluated downstream tasks here. Please see appendix B for details. Referring Expression Retrieval To assess the proposed method’s ability to capture localized properties, we first evaluate it for referring expression classification. Given an image and a set of masks, we generate an embedding for each mask and match the generated region embeddings to a set of text queries (referring expressions) encoded with the respective text encoder. The query mask whose localized embedding exhibits the highest cosine similarity with the text embedding of a referring expression is selected. We employ standard referring expression datasets: PhraseCut (Wu et al., 2020), RefCOCO, and RefCOCO+ (Kazemzadeh et al., 2014), reporting top-1, top-5, top-10 accuracy, mean Intersection over Union (mIoU) and overall Intersection over Union (oIoU).
@@
Class Retrieval Zero-shot classification requires classifying an image by matching its visual embedding to the closes textual description of all classes present in the dataset. Here, we classify a specific region of the image, indicated by a query mask on an object, leveraging two semantic segmentation datasets, PascalVOC (Everingham et al., 2015) and PascalContext (Mottaghi et al., 2014), and one instance segmentation dataset, MSCOCO (Lin et al., 2014). The performance is evaluated using the top-1, top-5, and top-10 accuracy. Finally, we evaluate the proposed method in a large-scale open-vocabulary setting on a subset of the OpenImagesV7 (Benenson and Ferrari, 2022), which offers mask annotations for a diverse array of objects across 350 unique classes.
@@
Localized Captioning Traditionally, image captioning models generate captions for entire images based on the visual representation provided by an image encoder. In contrast, we aim to evaluate MaskInversion’s ability to focus the captioner on a specific region, while maintaining contextual relevance. To this end, we leverage a pretrained image captioner, CLIPCap (Mokady et al., 2021), and provide it with the localized embedding token of a query mask to generate a caption. CLIPCap is trained on top of the CLIP vision encoder and feeds its [CLS] token to GPT-2(Radford et al., 2019) to produce a caption. Here, we feed the localized embeddings of MaskInversion as a drop-in replacement of CLIP’s [CLS] token to the captioner without finetuning. As no dataset directly supports this kind of evaluation, we adapted an existing dataset, PhraseCut(Wu et al., 2020). To quantitatively evaluate the generated localized captions, we match them to the set of ground-truth referring expressions for this image using the text encoder from CLIP (ViT-L/14 by OpenAI). We consider a caption as correct if its cosine similarity to the ground-truth referring expression for this mask is the highest among other referring expressions. The reported metric for this task is the top-1 accuracy.
4.2 Setup
@@
The proposed method is evaluated using pretrained CLIP vision-language models. For ViT-B/32, ViT-B/16, and ViT-L/14, we used the original weights from OpenAI (Radford et al., 2021), and for ViT-H/14, we used the weights "laion2b_s32b_b79k" from the OpenCLIP library (Cherti et al., 2023; Schuhmann et al., 2022). For the MaskInversion optimization, we use the AdamW optimizer(Kingma, 2014) with $10$ gradient descent iterations. We set the regularization parameter $\alpha$ to $5$ for RefCOCO and RefCOCO+, and to 0 for all other datasets.
@@
RedCircle ViT-B/16 21.5 72.3 90.3 42.5 43.2 32.7 42.5 43.3 33.5
FGVP ViT-B/16 35.9 83.5 95.2 42.6 43.2 33.3 48.0 48.7 38.0
AlphaCLIP ViT-B/16 34.0 80.0 93.6 43.4 44.0 38.1 44.2 44.7 39.7
@@
Table 1: Evaluation of MaskInversion on Referring Expression Retrieval. Given a query mask, the task is to retrieve the corresponding expression. $\ddagger$} indicates deviating evaluation settings where a pretrained region proposal is used and the prediction is counted as a hit if the matched region has an IoU>0.5>0.5; in this setting, several proposals could result in a hit. indicates reproduced results.
4.3 Comparison to the State-of-the-Art (SOTA)
Referring Expression Retrieval Table 1 presents the results on referring expression datasets. For some related approaches (CPT, GradCAM, ReCLIP, FGVP and RedCircle), as there is no directly comparable setting, we provide both the results as reported in their paper, and our reproduced results. Note that the original evaluation settings can vary for different methods. For reproduced results indicated by we adapt the evaluation setting to the case where ground-truth masks are used as described in Sec. 4.1. We used the code provided by the authors of each method, encode each image together with the ground-truth masks of MSCOCO, and match the resulting representation to the text embedding produced by the respective backbone. We further compare with the following baselines: CLIP refers to the general CLIP baseline by using the image CLS token, Crop uses the CLS token of cropped region by forwarding only this region through CLIP, and Masked Crop refers to forwarding the full image, but keeping only the masked region and replacing all other pixels with the average pixel value of the dataset. On the PhraseCut dataset, MaskInversion outperforms all baselines, regardless of the model size. On the RefCOCO and RefCOCO+ datasets, MaskInversion also achieves SOTA performance. In addition, MaskInversion’s performance scales well with the backbone size, establishing a new SOTA on every dataset when ViT-H/14 is used.
@@ 4.3 Comparison to the State-of-the-Art (SOTA)
RedCircle* 47.5 92.9 97.7 21.3 45.0 57.4 28.8 63.0 77.3 40.5 75.8 84.5
AlphaCLIP* 52.6 85.9 93.8 27.7 60.9 75.1 30.9 55.9 70.3 43.0 77.4 84.3
FGVP* 71.8 93.6 98.3 32.6 58.9 72.4 35.9 62.2 72.6 39.4 75.6 84.6
RIS* 78.0 95.2 98.1 38.1 62.7 74.3 43.6 65.3 72.4 34.5 66.5 75.8
B/32 MaskInversion 79.5 96.4 98.8 46.7 74.9 84.6 38.0 65.8 78.4 42.6 78.8 86.6
B/16 MaskInversion 85.4 96.4 98.8 58.1 83.7 90.5 44.7 71.6 83.0 46.3 80.4 87.9
@@
Class Retrieval Table 2 compares MaskInversion to other methods for the case of zero-shot class retrieval, keeping the same setting as for Referring Expression Retrieval. MaskInversion outperforms all other methods on all datasets on semantic segmentation datasets, such as PascalVOC and PascalContext. Furthermore, MaskInversion also exhibits good performance on the instance segmentation dataset COCO. These results demonstrate that MaskInversion can effectively direct the attention of the foundation model to multiple instances of the same object class at the same time, as well as to a single instance. Here, MaskInversion also outperforms the recently proposed AlphaCLIP (Sun et al., 2024), which fine-tunes CLIP on millions of mask-text pairs annotations, thereby demonstrating its ability to excel without the need to fine-tune CLIP. Finally, on OpenImagesV7, which features a much larger vocabulary of 350 classes, we can see that methods like AlphaCLIP perform well, as they are specifically trained for such tasks. Nonetheless, MaskInversion again outperforms all other methods we compared, demonstrating its capability to handle large vocabularies.
Method Backbone VOC Context COCO PhraseCut RefCOCO RefCOCO+ OpenImagesV7 Avg
MaskCLIP B/16 74.9 43.0 40.2 53.9 49.3 52.6 45.6 51.4
CLIPSurgery B/16 70.8 53.5 41.7 52.5 48.9 52.0 49.5 52.7
SCLIP B/16 64.3 43.0 33.4 37.2 40.7 42.4 45.5 43.8
Ours B/16 85.4 58.1 44.7 57.2 56.1 58.3 46.3 58.0
MaskCLIP L/14 55.1 33.2 29.3 47.6 43.2 47.2 32.5 41.2
CLIPSurgery L/14 78.3 46.4 47.7 47.2 47.3 50.9 45.5 51.9
SCLIP L/14 43.0 24.9 25.9 19.0 32.8 32.5 38.3 30.9
Ours L/14 91.0 59.0 56.0 60.2 56.1 60.2 48.7 61.6
MaskCLIP H/14 61.8 37.8 30.9 45.9 34.6 39.6 36.9 41.1
CLIPSurgery H/14 68.0 40.8 40.1 41.5 43.2 46.7 45.8 46.6
SCLIP H/14 38.2 20.7 19.8 15.2 20.7 20.7 35.6 24.4
Ours H/14 93.5 61.8 63.7 64.0 61.2 65.0 51.2 65.8
Table 3: Performance comparison of MaskInversion against different training-free method for different tasks. The reported numbers are Accuracies (higher is better).
Comparison with Training-free Methods We compare MaskInversion with recent training-free approaches for mask-conditioned CLIP representations: MaskCLIP Zhou et al. (2022), CLIPSurgery Li et al. (2023), and SCLIP Wang et al. (2024). These methods compute patch token representations and aggregate tokens within the mask region through average pooling to obtain localized embeddings. Using their official implementations, we evaluate these approaches across our benchmark suite while maintaining our evaluation protocol. Table 3 presents the comparative results across different CLIP architectures (ViT-B/16, ViT-L/14, and ViT-H/14). MaskInversion consistently outperforms training-free methods across nearly all datasets and model scales. The performance gap becomes particularly pronounced with larger backbone architectures, where MaskInversion demonstrates substantial improvements. For instance, with ViT-H/14, our method achieves gains of +31.7%, +21.0%, and +32.8% on VOC, Context, and COCO datasets respectively compared to the best performing baseline. Notably, while training-free methods show competitive performance with ViT-B/16, their effectiveness diminishes with larger architectures. In contrast, MaskInversion effectively leverages the increased capacity of larger models, showing consistent performance improvements across all scales. This suggests that our approach better utilizes the representational power of advanced CLIP models while maintaining robust performance across different architectural scales.
@@
Table 4: Mask Quality Ablation: Class retrieval accuracy on MSCOCO for mask variations, as well as for bounding boxes and SAM masks based on bounding boxes.
[…]
@@
Table 5: Gradient Decomposition: Runtime with and w/o gradient decomposition for different number of masks.
Method Acc
CLIP 20.1
AlphaCLIP 31.8
MaskInversion 48.4
Table 6: Localized captioning Ablation: We use CLIPCap to generate captions based on embeddings generated by CLIP, AlphaCLIP, and MaskInversion corresponding to the region highlighted by the mask. We report top-1 image-to-text retrieval accuracy.
4.4 Ablations